OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value if x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (gorv):

put x=3

OpenStudy (gorv):

f(3)=3-4=-1 so f(3) represent value of x-3 when x=3

OpenStudy (gorv):

first option @Jimboslice

OpenStudy (anonymous):

Thank you @gorv

OpenStudy (gorv):

:))

OpenStudy (anonymous):

@gorv will you help on another ?

OpenStudy (gorv):

yep just fire it buddy

OpenStudy (anonymous):

Brenda deposited some money in a bank. The bank paid her a simple interest every year. After t years, the amount of money f(t), in dollars, that the bank paid Brenda was $2000 greater than 10 times the number of years. Which statement best describes the steps to sketch a graph showing the relationship between f(t) and t? Label t on the x axis and f(t) on the y axis and plot a graph having slope = 2000 and y intercept = 10. Label t on the x axis and f(t) on the y axis and plot a graph having slope = 10 and y intercept = 2000. Label f(t) on the x axis and t on the y axis and plot a graph having slope = 10 and y intercept = 200. Label f(t) on the x axis and t on the y axis and plot a graph having slope = 2000 and y intercept = 200.

OpenStudy (gorv):

ok so time on x axis and f(t) on y axis f(t) = total amount=10t+2000 f(t)=10t+2000

OpenStudy (gorv):

slope =10 and intercept = 2000

OpenStudy (gorv):

option B

OpenStudy (anonymous):

Much appreciated @gorv